OpenStudy (anonymous):

Solve the system of equations x + y = 5 x - y = -9

OpenStudy (anonymous):

\[x + y = 5\]\[x - y = -9\]Add:\[2x=-4\] \[x=-2\]\[y=7\]

OpenStudy (anonymous):

x + y = 5 x - y = -9 add the above 2 equations to get 2x=-4 solve for x. we get x=-2 put this value in any one of the given equations and get value of y x + y = 5 -2+y=5 y=7 so x=-2 and y=7